OpenStudy (anonymous):

Cinema A sells a 12 pack of movie passes for 66 dollars. Cinema B sells a 5 pack of movie passess for 28 dollars. What is the cose per ticket for each cinema?

OpenStudy (anonymous):

$5.5 and $5.6

OpenStudy (gw2011):

If you look at the words "cost per ticket," this will tell you how to solve the problem. The word "per" means to divide.

OpenStudy (texaschic101):

Proportion time .... 12/66 = 1/x (cross multiply) 12 for 66 = 1 for x dollars 12x = 66 12x/12 = 66/12 x = 5.5 Cinema A : each ticket costs $ 5.50 5/28 = 1/x (cross multiply) 5 for 28 = 1 for x dollars 5x = 28 5x/5 = 28/5 x = 5.6 Cinema B : each ticket costs $ 5.60